OpenStudy (anonymous):

I need help with solving an equation for indicated variables, if someone could help me and check my work, that would be awesome!

OpenStudy (anonymous):

#1 \[p = \frac{ w }{ t } \] for w #2 \[P = \frac{ w }{ t } \] for t

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

There are more complicated ones than this >.>

OpenStudy (anonymous):

#3 \[E = \frac{ 9R }{ I }\]

OpenStudy (anonymous):

for I

OpenStudy (alekos):

what do you get for 1 and 2?

OpenStudy (anonymous):

tp = w wp = t >.>

OpenStudy (wolf1728):

1 is correct 2 is not

OpenStudy (anonymous):

Yeah I'm so confused with these >.<

OpenStudy (anonymous):

I know I have to get the indicated variable alone, but I don't know how >.<

OpenStudy (alekos):

t = w/p

OpenStudy (anonymous):

So those are the answers, but I want to know how to solve these! They're so hard but I want to learn them

OpenStudy (wolf1728):

when you have an equation such as a = b/c OR a/b = c/d you can do something that is called cross-multiplying

OpenStudy (alekos):

if p = w/t then if we multiply both sides by t we get pt = w then we divide both sides by p and we get t = w/p

OpenStudy (anonymous):

The second one is w = t/p

OpenStudy (anonymous):

Is that correct?

OpenStudy (alekos):

no w = pt If p = w/t then we multiply both sides by t and we get w=pt

OpenStudy (alekos):

do you follow?

OpenStudy (anonymous):

Oh ok

OpenStudy (alekos):

so what's your guess for no. 3

OpenStudy (anonymous):

9RE = I

OpenStudy (anonymous):

Oh and btw I have to state a property or explain each step for how I solved #3

OpenStudy (anonymous):

each step

OpenStudy (alekos):

If E=9R/I then we multiply both sides by I and we get EI = 9R Then we divide both sides by E and we get I = 9R/E

OpenStudy (alekos):

do you get it now?

OpenStudy (anonymous):

Ok, yeah #4 ak - r = on For k I got k = on/a + r

OpenStudy (anonymous):

Hard, but starting to make a lil sense

OpenStudy (anonymous):

I did #4 like 20min ago, so it may or may not be correct

OpenStudy (alekos):

First you add r to both sides

OpenStudy (anonymous):

ak = on + r

OpenStudy (anonymous):

Then isolate k right?

OpenStudy (alekos):

that's it. then you divide both sides by a

OpenStudy (anonymous):

\[k = \frac{ on + r }{ a }\] ???

OpenStudy (alekos):

beautiful!! you do get it!

OpenStudy (anonymous):

;3

OpenStudy (anonymous):

I just have 5 more of these to do xD

OpenStudy (alekos):

ok

OpenStudy (anonymous):

Which shows the equation \[c = a \left(\begin{matrix}w \\ 150\end{matrix}\right)\] correctly solved for the variable w? A. w = a(150c) B. w = a(c/150) C. w = 150/ac D. w = 150c/a

OpenStudy (anonymous):

D!

OpenStudy (alekos):

Yes it's D

OpenStudy (anonymous):

#35 w = gm Solve for m #36 Q = 1/2P + 15 Solve for P #37 I = V/R Solve for R #38 y = mx + b Solve for m

OpenStudy (anonymous):

Let me do those real quick ;3

OpenStudy (anonymous):

#35 w/g = m #36 Tricky one \[\frac{ 2 }{ 1 } \times Q - 15 = p \] #37 IV = R #38 \[\frac{ y-b }{ x } = m\]

OpenStudy (alekos):

#35 is correct

OpenStudy (alekos):

#36 is close. P = 2(Q-15)

OpenStudy (anonymous):

Oh just simplify the 2/1 to a 2?

OpenStudy (alekos):

#37 R = V/I

OpenStudy (anonymous):

my teacher is really weird about fractions, sometimes he just says to leave them in that form and he knows you understand it. So I think he might take 2/1

OpenStudy (alekos):

and #38 is right!!

OpenStudy (anonymous):

But simplifying it does make it more clear c;

OpenStudy (alekos):

that's fine. for #36 you can go P=(2/1) x (Q-15) but it's normally expressed as I first wrote it

OpenStudy (alekos):

well done!!

OpenStudy (alekos):

i hope you have a better understanding of algebraic manipulation

OpenStudy (anonymous):

Ok I changed #36, it was a simple error and I understand it now :) Thanks so much, I wish I could give you a million medals >.>

OpenStudy (anonymous):

You've been here with me for like a whole hour :P

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

If you don't mind me asking, what grade are you?

OpenStudy (alekos):

I'm a high school tutor

OpenStudy (anonymous):

Cool ;3

OpenStudy (anonymous):

I'm solving this problem now \[2.80 = \frac{ 9(70) }{ I }\]

OpenStudy (anonymous):

In baseball, the equation \[E = \frac{ 9R }{ I }\] gives a pitchers run average E, where R is the number of earned runs the player allowed and I is the number of innings pitched. Last season, a pitcher had earned run average of 2.80 and allowed 70 earned runs. how many innings did the pitcher pitch last season?